carmelle:

If 0 < f ≤ 90 and cos(22f − 1) = sin(7f + 4), what is the value of f? A. f = 3 B. f = 4 C. f = 5 D. f = 6

Vocaloid:

cos(theta) = sin(90-theta) so cos(22f-1) = sin( 90 - [22f - 1]) = sin(7f + 4) since you now have sin on both sides you can set up 90 - (22f-1) = 7f + 4 and solve from there

carmelle:

ah ok, I got f = 3 as my final answer. thank you!